# New Advances in Phonons: From Band Topology to Quasiparticle Chirality

### v1

@@ V From Topological Phonons to Chiral Phonons
In this section, we will review several new fundamental concepts related to the chirality of circularly polarized phonons, including their relationships and corresponding experimental benchmarks illustrated with material examples. Additionally, we will review chiral phonon-driven physical phenomena and explore their potential applications in various fields.
@@
Figure 32: The previous definition of chiral phonons based on angular momentum is convention-dependent and not well-defined, as angular momentum is a pseudo-vector. To address this issue, an alternative definition based on phonon helicity is proposed Zhang et al. (2025a). Phonon helicity is a pseudo-scalar and thus convention-independent and well-defined. The relationships between angular momentum, helicity, and reprensentive atomic motions are illustrated as follows: (a) Phonon modes with zero helicity can exhibit three possibilities. The first two cases represent nonchiral phonons under the helicity-based definition, whereas they would be classified as chiral phonons under the angular momentum-based definition. (b) Phonon modes with nonzero helicity must have nonzero angular momentum, and the sign of the helicity directly indicates the chirality of the circularly polaried phonon modes. Adapted from Ref. Zhang et al. (2025a).
V.1 Basic Concepts Related to Chiral Phonons
@@ V.1.1 Phonon Angular Momentum
where $\bm{u}_{\alpha}(R_{l},\tau_{\kappa})$ represents the displacement of the $\kappa$-th atom in the $l$-th unit cell along $\alpha$ direction. With this, one can define phonon AM, i.e., $l_{\alpha,\nu\bm{q}}$, for a specific phonon mode $\bm{\epsilon}_{\nu\bm{q}}$ at momentum $\bm{q}$ in the equilibrium state as McLellan (1988); Zhang and Niu (2014); Kishine et al. (2020); Zhang et al. (2025a):
@@
From a semi-classical perspective, $\bm{l}_{\nu\bm{q}}$ can be interpreted as the summation of the circular polarization of atoms in the primitive unit cell. Consequently, a phonon mode with $l_{\alpha,\nu\bm{q}}\neq 0$ corresponds to atomic motions where the net circular polarization of all atoms in the unit cell is nonzero. The total AM of the system, denoted as $J^{\textrm{ph}}$, is defined as the ensemble average of the AM of each phonon mode across the entire Brillouin zone. Mathematically, this can be expressed as: $J^{\textrm{ph}}=\sum_{\alpha,\nu,\bm{q}}\langle l_{\alpha,\nu\bm{q}}\rangle$, where $\langle l_{\alpha,\nu\bm{q}}\rangle$ represents the average angular momentum of the phonon mode $\nu$ at momentum $\bm{q}$. This framework provides a comprehensive understanding of the phonon Edelstein effect and phonon angular momentum in crystalline materials Xu et al. (2018); Hamada et al. (2018); Hamada and Murakami (2020b); Zhang et al. (2024a).
@@ V.1.2 Phonon Helicity
Previously, chiral phonons (circularly polarized phonons), were referred to as phonon modes with nonzero AM and were extensively studied in 2D systems. However, chirality inherently requires three degrees of freedom: two for circular motion and one for the propagation direction. Additionally, AM is a pseudovector, making its definition convention-dependent under SO(3) operations. For example, the algebraic sign of the AM undergoes inversion when adopting left-handed versus right-handed coordinate system conventions. This limitation underscores the need for a more robust and convention-independent definition of chiral phonons, such as one based on phonon helicity Zhang et al. (2025a).
@@
For phonon modes with zero helicity, the phonon AM can either be perpendicular to $\bm{q}$ or zero, as illustrated in Fig. 32 (a). Consequently, phonon modes defined as chiral based on nonzero AM can be achiral under the definition of phonon helicity. This ambiguity highlights that chiral phonons are not well-defined in purely 2D systems, as the propagation direction will always be perpendicular to the AM. In previous studies, to address the lack of a third degree of freedom in 2D systems, the $z+$ direction was often assumed as the propagation direction to define the chirality of phonons. In 3D systems, however, chiral phonons are well-defined as phonon modes with nonzero helicity, where the phonon AM is not perpendicular to $\bm{q}$. Thus, chiral phonons are characterized by nonzero helicity, which inherently implies nonzero AM, and the sign of the helicity reflects the chirality of the circularly polarized phonons, as shown in Fig. 32 (b). This definition provides a more robust and convention-independent framework for identifying and studying chiral phonons in 3D systems, and it is also closely related to the chiral charge density waves Luo and Dai (2023); Zhang et al. (2024b); Yang et al. (2024a); Romao and Juraschek (2024).
@@ V.1.3 Phonon Pseudo-angular Momentum
Because neither the phonon AM nor the helicity is easily observable in experiments, the pseudo-angular momentum (PAM) is introduced as a practical benchmark for identifying chiral phonons Yao et al. (2008); Zhang and Niu (2015); Zhang and Murakami (2022); Zhang et al. (2025a); Bourgeois et al. (2025); Zhang et al. (2025b, 2023b); Zhu et al. (2018); Ishito et al. (2023a, b); Komiyama et al. (2022). For a phonon mode ${u}_{\kappa\alpha,\bm{q}}$ at momentum $\bm{q}$ with little group $\mathcal{C}_{n}$, the representation matrix of $\mathcal{C}_{n}$ can be derived as follows:
@@
$P_{\kappa^{\prime}\kappa}$ is a permutation matrix describing the transformation between the sublattice $\kappa$ and $\kappa^{\prime}$. $\bm{G}$ $=\sum_{i=1}^{3}n_{i}\bm{b}_{i}$, where $\bm{b}_{i}$ is the reciprocal lattice vector and $n_{i}$ is an integer. The matrix $P_{\kappa^{\prime}\kappa}$ is a permutation matrix that describes the transformation between the sublattices $\kappa$ and $\kappa^{\prime}$ under the symmetry operation. Meanwhile, $\bm{G}$ is a reciprocal lattice vector defined as:
@@
$l_{rot}$ represents the pure rotational part of the PAM, corresponding to the eigenvalue of the rotational symmetry $\mathcal{C}_{n}$. $l_{\bm{\tau}}=\frac{\bm{q}\cdot\bm{\tau}_{m/n}}{2\pi/n}$ accounts for the contribution from the fractional translation $\bm{\tau}_{m/n}$ along the rotational axis. This decomposition reveals that the PAM $l_{ph}$ becomes $\bm{q}$-dependent in systems with screw rotational symmetries $\mathcal{C}_{n,\bm{\tau}_{m/n}}$. As a result, $l_{ph}$ can take non-integer values, in contrast to systems with pure rotational symmetries where it is quantized to integers modulo $n$ Zhang and Murakami (2022); Zhang et al. (2025a).
@@
Previous studies have used PAM as a benchmark for circularlly polarized phonons defined by AM, suggesting an intrinsic relationship between these two. However, PAM has no intrinsic connection to either phonon angular momentum or phonon helicity, meaning it cannot serve as an experimental benchmark for observing chiral phonons. Nevertheless, in certain cases, PAM can be used for identifying nonzero AM, combined with symmetry analysis, first-principles calculations, etc., as shown in Zhu et al. (2018); Ishito et al. (2023b); Zhang et al. (2023b, 2025a); Yang et al. (2024b); Che et al. (2024); Ishito et al. (2023a); Bourgeois et al. (2025); Zhang et al. (2025b). Zhang et al. (2025a) provides a comprehensive explanation of the distinction between PAM and AM from a definitional perspective. The rotation operator in the $O(3)$ group with an arbitrary angle $\phi$ along the ${\alpha}$-th direction can be expressed as:
@@
Since phonon AM is the expectation value of $\mathfrak{M}_{\alpha}$, it can be defined in any system, regardless of its symmetry. The definition in Eq. (81) demonstrates that AM solely encapsulates information about the phonon polarization and atomic motion within the primitive cell. It is independent of the relative phases between different atoms, reflecting the local property of the phonon mode, as discussed in Zhang et al. (2025a).
@@
Phonon PAM is defined by the eigenvalue of $\mathcal{C}_{n}$, as shown in Eq. (85), and thus can only be defined at $\mathcal{C}_{n}$-invariant momenta. Furthermore, PAM encodes information about the relative phase between $\mathcal{C}_{n}$-related atoms, potentially extending across multiple primitive cells. From this perspective, PAM can be regarded as a global property of a specific phonon mode Zhang et al. (2025a).
@@
In conclusion, phonon AM can be defined in any systems and reflects the local properties of a phonon mode. In contrast, PAM is defined only at (screw) rotation-invariant $\bm{q}$ points, as it reflects the eigenvalues of $\mathcal{C}_{n}$, revealing global information about the phonon mode. There is no intrinsic one-to-one relationship between AM and PAM. Therefore, to observe chiral phonons by detecting PAM, first-principles calculations and symmetry analysis are essential, as demonstrated in studies on Te Zhang et al. (2023b) and $\alpha-$HgS Ishito et al. (2023b). The selection rules for circularly polarized Raman scattering in systems with rotational symmetries are illustrated in Fig. 35.
@@
Figure 33: The connection and differences between topological phonons and specifically defined chiral phonons (circularly polarized phonons defined by AM) can be summarized as follows: Weyl phonons are both topological and chiral. Topological phonons are characterized by topological invariants, leading to rich classifications based on the type of invariant, such as Weyl phonons, Dirac phonons, and nodal-line phonons. When the topological invariant is nonzero, the phonon mode is topological, exhibiting features like topological surface states (“bulk-surface correspondence”) and unique pseudospin textures around the degeneracy points. Weyl phonons, for instance, are both topological and chiral, with a notable example being the twofold-degenerate quadruple Weyl phonon in BaPtGe. Chiral phonons, or the circularly polarized phonons, on the other hand, have been defined in various ways. Earlier studies defined them based on nonzero AM, but this definition is not well-defined. More recently, chiral phonons have been redefined using nonzero phonon helicity, which is a more robust and convention-independent criterion. When a phonon mode has nonzero helicity, atoms in real space exhibit circular vibrations, reflecting the chiral nature of the mode. Thus, while topological phonons are defined by their topological invariants, chiral phonons are defined by their helicity, with Weyl phonons serving as an example of modes that are both topological and chiral. Adapted from Ref. Zhang et al. (2025b).
@@
Figure 34: A group of band structures for a system along the $-A$-$\Gamma$-$A$ path, which preserves $\mathcal{C}_{3}$ rotational symmetry. Here, $p$ denotes the pseudo-angular momentum (PAM) value for each band, and $\chi$ represents the monopole charge of the Weyl phonons. It is important to note that the PAM for a specific phonon mode can be non-integer in systems with $\mathcal{C}_{n}$ screw rotational symmetry. Adapted from Ref. Tsirkin et al. (2017).
@@
Figure 35: The selection rules for circularly polarized Raman scattering in the Stokes [(a1-d1)] and anti-Stokes [(a2-d2)] process for the systems with $\mathcal{C}_{n}={2,3,4,6}$ rotation symmetries, where the incident and scattered light propagate along the rotation axis. Each arrow represents the eigenvalue of a phonon mode in the complex plane, corresponding to its PAM value. In each figure, phonon modes marked by the black arrow are Raman active in both the $\sigma^{+}/\sigma^{+}$ and $\sigma^{-}/\sigma^{-}$ processes, phonon modes marked by the blue (red) arrows are active in the $\sigma^{+}/\sigma^{-}$ ($\sigma^{-}/\sigma^{+}$) process. These selection rules highlight the dependence of Raman activity on the PAM values and the circular polarization of the incident and scattered light. Adapted from Ref. Zhang et al. (2025a).
V.3 Connections between Topological Phonon & Chiral Phonon
In this subsection, we will explore the intricate relationship and distinct characteristics of topological phonons and specifically defined chiral phonons, i.e., circularly polarized phonons defined by AM, as illustrated in Fig. 33. Our focus will be on Weyl phonons, which exhibit both topological and chiral properties, and we will provide a detailed analysis of their features from the definition point of view.
V.3.1 Weyl Phonons are Both Topological and Chiral

### v2

@@ V From Topological Chirality to Rotational Chirality
Since “chiral phonons” have different meanings in different research field, it is essential to clarify the underlying assumptions and conditions when discussing phonon chirality. In this review, we adopt “circularly polarized phonons” to describe phonon modes that carry nonzero angular momentum. This terminology is also intuitive and emphasizes the circular motion of atoms in real space. Below, we review fundamental concepts related to circularly polarized phonons and topologically chiral phonons, including their relationships and corresponding experimental benchmarks illustrated with material examples. Additionally, we will review circularly polarized phonon-driven physical phenomena and explore their potential applications in various fields.
@@
Figure 30: The previous definition of chiral phonons based on angular momentum is convention-dependent and not well-defined, as angular momentum is a pseudo-vector. To address this issue, an alternative definition based on phonon helicity is proposed zhang2025thechirality . Phonon helicity is a pseudo-scalar and thus convention-independent and well-defined. The relationships between angular momentum, helicity, and reprensentive atomic motions are illustrated as follows: (a) Phonon modes with zero helicity can exhibit three possibilities. The first two cases represent nonchiral phonons under the helicity-based definition, whereas they would be classified as chiral phonons under the angular momentum-based definition. (b) Phonon modes with nonzero helicity must have nonzero angular momentum, and the sign of the helicity directly indicates the chirality of the circularly polaried phonon modes. Adapted from Ref. zhang2025thechirality .
V.1 Basic Concepts Related to Circularly Polarized Phonons
@@ V.1.1 Phonon Angular Momentum
where $\bm{u}_{\alpha}(R_{l},\tau_{\kappa})$ represents the displacement of the $\kappa$-th atom in the $l$-th unit cell along $\alpha$ direction. With this, one can define phonon AM, i.e., $l_{\alpha,\nu\bm{q}}$, for a specific phonon mode $\bm{\epsilon}_{\nu\bm{q}}$ at momentum $\bm{q}$ in the equilibrium state as mclellan1988angular ; zhang2014angular ; kishine2020chirality ; zhang2025thechirality :
@@
From a semi-classical perspective, $\bm{l}_{\nu\bm{q}}$ can be interpreted as the summation of the circular polarization of atoms in the primitive unit cell. Consequently, a phonon mode with $l_{\alpha,\nu\bm{q}}\neq 0$ corresponds to atomic motions where the net circular polarization of all atoms in the unit cell is nonzero. The total AM of the system, denoted as $J^{\textrm{ph}}$, is defined as the ensemble average of the AM of each phonon mode across the entire Brillouin zone. Mathematically, this can be expressed as: $J^{\textrm{ph}}=\sum_{\alpha,\nu,\bm{q}}\langle l_{\alpha,\nu\bm{q}}\rangle$, where $\langle l_{\alpha,\nu\bm{q}}\rangle$ represents the average angular momentum of the phonon mode $\nu$ at momentum $\bm{q}$. This framework provides a comprehensive understanding of the phonon Edelstein effect and phonon angular momentum in crystalline materials xu2018inverse ; hamada2018phonon ; hamada2020phonon ; zhang2024observation .
@@ V.1.2 Phonon Helicity
Previously, circularly polarized phonons, were referred to as phonon modes with nonzero AM and were extensively studied in 2D systems. However, chirality inherently requires three degrees of freedom: two for circular motion and one for the propagation direction. Additionally, AM is a pseudovector, making its definition convention-dependent under SO(3) operations. For example, the algebraic sign of the AM undergoes inversion when adopting left-handed versus right-handed coordinate system conventions. This limitation underscores the need for a more robust and convention-independent definition of chiral phonons, such as one based on phonon helicity liu2022probing ; hu2021phonon ; zhang2025thechirality .
@@
For phonon modes with zero helicity, the phonon AM can either be perpendicular to $\bm{q}$ or zero, as illustrated in Fig. 30 (a). Consequently, phonon modes defined as chiral based on nonzero AM can be achiral under the definition of phonon helicity. This ambiguity highlights that chiral phonons are not well-defined in purely 2D systems, as the propagation direction will always be perpendicular to the AM. In previous studies, to address the lack of a third degree of freedom in 2D systems, the $z+$ direction was often assumed as the propagation direction to define the chirality of phonons. In 3D systems, however, chiral phonons are well-defined as phonon modes with nonzero helicity, where the phonon AM is not perpendicular to $\bm{q}$. Thus, chiral phonons are characterized by nonzero helicity, which inherently implies nonzero AM, and the sign of the helicity reflects the chirality of the circularly polarized phonons, as shown in Fig. 30 (b). This definition provides a more robust and convention-independent framework for identifying and studying chiral phonons in 3D systems, and it is also closely related to the chiral charge density waves luo2023transverse ; zhang2024understanding ; yang2024incommensurate ; romao2024phonon .
@@ V.1.3 Phonon Pseudo-angular Momentum
Since it requires considerable effort to directly observe phonon angular momentum zhang2024observation , the pseudo-angular momentum (PAM) has been introduced as a practical benchmark for identifying circularly polarized phonons yao2008valley ; zhang2015chiral ; zhang2022chiral ; zhang2025thechirality ; bourgeois2025strategy ; zhang2025weyl ; zhang2023weyl ; zhu2018observation ; ishito2023chiral ; ishito2023truly ; komiyama2022physics . However, we note that this approach is applicable only in limited cases. For a phonon mode ${u}_{\kappa\alpha,\bm{q}}$ at momentum $\bm{q}$ with little group $\mathcal{C}_{n}$, the representation matrix of $\mathcal{C}_{n}$ can be derived as follows:
@@
$P_{\kappa^{\prime}\kappa}$ is a permutation matrix describing the transformation between the sublattice $\kappa$ and $\kappa^{\prime}$. $\bm{G}$ $\sum_{i=1}^{3}n_{i}\bm{b}_{i}$, where $\bm{b}_{i}$ is the reciprocal lattice vector and $n_{i}$ is an integer. The matrix $P_{\kappa^{\prime}\kappa}$ is a permutation matrix that describes the transformation between the sublattices $\kappa$ and $\kappa^{\prime}$ under the symmetry operation. Meanwhile, $\bm{G}$ is a reciprocal lattice vector defined as:
@@
$l_{rot}$ represents the pure rotational part of the PAM, corresponding to the eigenvalue of the rotational symmetry $\mathcal{C}_{n}$. $l_{\bm{\tau}}=\frac{\bm{q}\cdot\bm{\tau}_{m/n}}{2\pi/n}$ accounts for the contribution from the fractional translation $\bm{\tau}_{m/n}$ along the rotational axis. This decomposition reveals that the PAM $l_{ph}$ becomes $\bm{q}$-dependent in systems with screw rotational symmetries $\mathcal{C}_{n,\bm{\tau}_{m/n}}$. As a result, $l_{ph}$ can take non-integer values, in contrast to systems with pure rotational symmetries where it is quantized to integers modulo $n$ zhang2022chiral ; zhang2025thechirality .
@@
Previous studies have used PAM as a benchmark for circularlly polarized phonons defined by AM, suggesting an intrinsic relationship between these two. However, PAM has no intrinsic connection to either phonon angular momentum or phonon helicity, meaning it cannot serve as an experimental benchmark for observing circularly polarized phonons. Nevertheless, in certain cases, PAM can be used for identifying nonzero AM, combined with symmetry analysis, first-principles calculations, etc., as shown in zhu2018observation ; ishito2023truly ; zhang2023weyl ; zhang2025thechirality ; yang2025inherent ; che2025magnetic ; ishito2023chiral ; bourgeois2025strategy ; zhang2025weyl . zhang2025thechirality ; streib2021difference provides a comprehensive explanation of the distinction between PAM and AM from a definitional perspective. The rotation operator in the $O(3)$ group with an arbitrary angle $\phi$ along the $\alpha$}-th direction can be expressed as:
@@
Since phonon AM is the expectation value of $\mathfrak{M}_{\alpha}$, it can be defined in any system, regardless of its symmetry. The definition in Eq. (89) demonstrates that AM solely encapsulates information about the phonon polarization and atomic motion within the primitive cell. It is independent of the relative phases between different atoms, reflecting the local property of the phonon mode, as discussed in zhang2025thechirality .
@@
Phonon PAM is defined by the eigenvalue of $\mathcal{C}_{n}$, as shown in Eq. (93), and thus can only be defined at $\mathcal{C}_{n}$-invariant momenta. Furthermore, PAM encodes information about the relative phase between $\mathcal{C}_{n}$-related atoms, potentially extending across multiple primitive cells. From this perspective, PAM can be regarded as a global property of a specific phonon mode zhang2025thechirality .
@@
In conclusion, phonon AM can be defined in any systems and reflects the local properties of a phonon mode. In contrast, PAM is defined only at (screw) rotation-invariant $\bm{q}$ points, as it reflects the eigenvalues of $\mathcal{C}_{n}$, revealing global information about the phonon mode. There is no intrinsic one-to-one relationship between AM and PAM. Therefore, to observe circularly polarized phonons by detecting PAM, first-principles calculations and symmetry analysis are essential, as demonstrated in studies on Te zhang2023weyl and $\alpha$-HgS ishito2023truly . The selection rules for circularly polarized Raman scattering in systems with rotational symmetries are illustrated in Fig. 33.
@@
Figure 31: Schematic diagram showing the interconnection between chirality, topology, and Berry curvature for phonons, using Weyl phonons as a key example at the overlap of all three properties. Following the terminology of this review, the “chiral phonon” in the center of the right circle should be labeled a “circularly polarized phonon”.Topological phonons are characterized by topological invariants, leading to rich classifications based on the type of invariant, such as Weyl phonons, Dirac phonons, and nodal-line phonons. When the topological invariant is nonzero, the phonon mode is topological, exhibiting features like topological surface states (“bulk-surface correspondence”) and unique pseudospin textures around the degeneracy points. Weyl phonons, for instance, are both topological and chiral, with a notable example being the twofold-degenerate quadruple Weyl phonon in BaPtGe. Chiral phonons, or the circularly polarized phonons, on the other hand, have been defined in various ways. Earlier studies defined them based on nonzero AM, but this definition is not well-defined. More recently, chiral phonons have been redefined using nonzero phonon helicity, which is a more robust and convention-independent criterion. When a phonon mode has nonzero helicity, atoms in real space exhibit circular vibrations, reflecting the chiral nature of the mode. Thus, while topological phonons are defined by their topological invariants, chiral phonons are defined by their helicity, with Weyl phonons serving as an example of modes that are both topological and chiral. Adapted from Ref. zhang2025weyl .
@@
Figure 32: A group of band structures for a system along the $\Gamma-AA$ path, which preserves $\mathcal{C}_{3}$ rotational symmetry. Here, $p$ denotes the pseudo-angular momentum (PAM) value for each band, and $\chi$ represents the monopole charge of the Weyl phonons. It is important to note that the PAM for a specific phonon mode can be non-integer in systems with $\mathcal{C}_{n}$ screw rotational symmetry. Adapted from Ref. tsirkin2017composite .
@@
Figure 33: The selection rules for circularly polarized Raman scattering in the Stokes [(a1-d1)] and anti-Stokes [(a2-d2)] process for the systems with $\mathcal{C}_{n}={2,3,4,6}$ rotation symmetries, where the incident and scattered light propagate along the rotation axis. Each arrow represents the eigenvalue of a phonon mode in the complex plane, corresponding to its PAM value. In each figure, phonon modes marked by the black arrow are Raman active in both the $\sigma^{+}/\sigma^{+}$ and $\sigma^{-}/\sigma^{-}$ processes, phonon modes marked by the blue (red) arrows are active in the $\sigma^{+}/\sigma^{-}$ ($\sigma^{-}/\sigma^{+}$) process. These selection rules highlight the dependence of Raman activity on the PAM values and the circular polarization of the incident and scattered light. Adapted from Ref. zhang2025thechirality .
V.3 Connections between Topological Phonon & Circularly Polarized Phonons
This subsection examines the relationship between topological phonons and circularly polarized phonons characterized by nonzero angular momentum (Fig. 31). We focus on Weyl phonons, which exhibit both topological and circularly polarized properties, analyzing their defining features. Moreover, such systems often host electronic topology (e.g., chiral fermions), and a growing body of research points to a connection between this electronic topology and phonon chirality, a detailed examination of which lies beyond the scope of this review cui2023chirality ; luo2023evidence ; CoSnS_CP_2025 ; Che2025May ; zhang2025electronic .
V.3.1 Weyl Phonons are Both Topological and Circularly Polarized